# DualMAR: Medical-Augmented Representation from Dual-Expertise Perspectives

### v1

@@ 3.2 Generate “Knowledge Scholar”
3.2.1 Node and Edge Normalization.
Our approach for normalizing the combined KG $\mathcal{G}_{H}$ involves two key steps: (1) Uniform encoding system for entities: In addition to phenotype nodes sourced from HPO, we align disease and drug entities using the ICD-9-CM Organization et al. (1988) and ATC Nahler and Nahler (2009) coding systems, both of which are widely used in EHR data. A cross-referencing method is employed to convert $\mathcal{G}_{M}$ and $\mathcal{G}_{N}$ into a shared hierarchy. Given that drugs are not the primary focus, we use ATC-4 to simplify the numerous synergistic interactions in DrugBank. (2) Clustering duplicated or unmatched triples: The LLM-prompting process is run multiple times, which can generate duplicated nodes and edges. Moreover, cross-referencing files may not cover all concepts, and unmatched entities leave noise in KG. To address these problems, we calculate cosine similarity for nodes and edges based on their embeddings. Using a predefined threshold $\theta$, we group entities within categories such as “disease”, “drug”, “phenotype”, and “other”, and apply hierarchical clustering to reorganize edges.
Table 1: Statistical Comparison among KGs.
(* means results might fluctuate by different parameter settings.)
@@
3.2.2 Advancement Analysis
To demonstrate the advancements of our proposed Diagnosis KG, we compare it with two recent health-related KGs: GraphCare Jiang et al. (2023) and PrimeKG Chandak et al. (2023). A statistical summary of KGs is presented in Table 1. The analysis highlights the following advantages of our KG:
@@ 3.2.2 Advancement Analysis
    Compared to GraphCare, our KG is larger in scale and built on the widely-used ICD-9-CM system, offering hierarchical and precise knowledge for predictive healthcare. Additionally, while GraphCare relies on zero-shot generation and subgraph sampling, our KG provides more stable and accurate results.
@@
    Compared to PrimeKG, our KG includes more edge types, primarily augmented by LLM-generated triples. The LLM reasoning process not only enhances connectivity within the KG but also converts unstructured text into structured triples, making information retrieval more efficient from single-modal data.
Although PrimeKG covers a broader range of medical concepts, it is more prone to data quality issues that can introduce bias into EHR predictions. For example, genes and proteins included in PrimeKG are seldom considered in physicians’ decision-making processes, particularly when it comes to diagnosis. In contrast, our diagnosis-driven KG contains numerous normalized triples with diverse edges, enhanced by LLMs’ in-context learning capabilities, making it highly reproducible and well-suited for various diagnosis-related healthcare predictions.
@@ 3.3 Polar-space KG Embedding
Our generated KG features a bi-hierarchical structure, with diagnostic and phenotype nodes organized across multiple levels. Traditional KG embedding methods often focus primarily on contextual information, but our framework gives equal importance to both contextual and hierarchical information. We employ a polar-space embedding Zhang et al. (2020) method upon the Euclidean space due to its superiority in efficiently capturing hierarchical structures compared to other Euclidean embedding. While non-Euclidean methods are also designed for hierarchical structures, they primarily focus on hierarchy alone. In contrast, polar-space embedding not only capture both hierarchical and contextual information simultaneously but also provide a more computationally efficient and interpretable framework, making it well-suited for integrating complex medical knowledge into predictive models. A comparison among KG embedding methods Bordes et al. (2013); Nickel and Kiela (2017); Sun et al. (2019); Fionda and Pirrò (2020) is provided in Appendix B.
To describe medical concepts from both contextual and hierarchical perspectives, we leverage two properties of the polar coordinate system: (1) Radial coordinate: map entities across different levels of the hierarchy; (2) Angular coordinate: map entities at the same level based on contextual information. We denote entities, including head $\mathbf{h}$ and tail $\mathbf{t}$, as $\mathbf{e}$, and relation embeddings as $\mathbf{r}$. Subscripts, such as $\mathbf{e}_{r}$ and $\mathbf{e}_{a}$, represent the modulus and phase components, respectively. Each entry of $\mathbf{h}_{r,a}$ or $\mathbf{t}_{r,a}$ corresponds to a radial or angular value, while each entry of $\mathbf{r}_{r,a}$ represents a scaling transformation between the corresponding head and tail. The radial and angular coordinates of the entity embeddings can then be formulated as follows:
@@
where $\|\cdot\|_{1}$, $\|\cdot\|_{2}$ denote the $\ell_{1}$, $\ell_{2}$ norm. The final circular arc distance $d(\mathbf{h},\mathbf{t})$ can be consider as the weighted sum of both radial and angular distances. Note that, we can get $\mathbf{h}$ and $\mathbf{t}$ by concatenating both radial and angular vectors. Such embedding model then be pretrained individually through link prediction tasks before integrating into the predictive model, which can help us understand priors from KG into trainable embeddings. The regular negative sampling loss function $L$ is adopted for optimization:
@@
where $\gamma$ and $\sigma$ are marginal value and sigmoid function respectively, and $({h^{\prime}},r,{t^{\prime}})$ is a negative triple. Therefore, we can get diagnosis embedding matrix $\mathbf{E_{\mathcal{D}}}$ for condition codes, which is the initial nodes features for “Local Expert”.
3.4 Continual Enhancement from “Local Expert”
The continual enhancement process unfolds in two steps: graph learning and proxy-task learning, which together refine the priors from the “Knowledge Scholar” by fully leveraging EHR data. The proposed encoder-decoder framework not only emphasizes the comprehensive understanding of medical patterns but also seamlessly integrates local knowledge into the learning process, further capturing critical medical patterns for precise diagnosis.
3.4.1 Encoder Module.
The backbone of the encoder module contains graph learning and attention, which adjusts code-level embeddings and progressively refines them into comprehensive patient representations.
@@ 3.4 Continual Enhancement from “Local Expert”
$\mathbf{W}_{c}$, $\mathbf{W}_{u}$, and $\mathbf{W}_{v}$ are weighted matrices. The attention score $\alpha_{\tau}$ measures the distribution of medical codes of admission, and $\beta$ measures the distribution across admissions. Note that, we project each admission $\mathbf{v}_{\tau}$ to $\tilde{\mathbf{v}}_{\tau}$ for fitting the patient dimension through a single layer with $\mathbf{W}_{u}$. Detailed explanation about attention is provided in Appendix C.
The proposed encoder module ensures that both the specificity of individual medical encounters and the comprehensive history of a patient’s health are reflected in the final patient embedding $\mathbf{p}$, laying a robust foundation for the subsequent self-supervised learning phase.
3.4.2 Proxy-Task Learning.
Given patient embedding $\mathbf{p}$, we introduce a novel proxy-task learning paradigm. Some predictive models in EHR focus on customized proxy tasks to utilize single admission which is discarded in training process. However, designed labels of proxy tasks should not only reflect historical pattern of previous admissions, but also closely connect to downstream labels for learned patient embeddings $\mathbf{p}$. Therefore, we harness the insight of lab results in our learning model, which is unexplored as auxiliary features within pretraining module. By pretraining the model on laboratory results, we aim to mimic the diagnostic reasoning process of real physicians, who often rely on lab data for accurate assessments.
We denotes lab results as $\mathcal{L}$, and the designed proxy task makes predictive model inherently understands intricate relations between abnormal lab tests and health conditions. Considering patterns always change across different lab tests, we further categorize lab results into distinct groups, which can fully cover involved test items in EHR data. Here we take lab tests in MIMIC as an example:
@@
    Hematology (# items: 417): Analyze blood components like red and white blood cells to monitor conditions.
@@
Hence, we categorize lab tests as $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ in order, and predict lab results $\hat{y}_{i,j}=\text{Decoder}(\mathbf{p})$ based on both single-admission and multi-admission data in training set, where $y_{i,j}=1$ means lab code $l_{j}$ shows abnormal results, and a multi-layer perceptron (MLP) is used as decoder for learning patterns in different lab categories. We only focus on abnormal lab tests across admissions, so the ground truth $[l_{1},l_{2},...,l_{|\mathcal{L}|}]$ are labeled as multi-hot vectors. To ensure the lab embeddings are well-aligned with each category, the proxy task is designed to be completed in two steps:
  • Joint Training. We can get preliminary trained parameters in encoder $\Theta_{e}$ and optimize prediction through the average binary cross entropy (BCE) loss functions across three categories, where we can adjust patient embedding $\tilde{\mathbf{p}}$ through updated parameters.
    $L_{j}=\frac{1}{|\mathcal{Y}|}\sum_{i=1}^{|\mathcal{Y}|}\text{BCE}(\mathbf{\hat{y}}_{i},\mathbf{y}_{i}),\>\mathcal{Y}=\{\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}\}$ (12)
    where $\mathbf{y}_{i}$ and $\mathbf{\hat{y}}_{i}$ are ground truth and prediction of lab results and $i$ means the $i$-th lab category.
@@ 4.1.1 Datasets
To evaluate our proposed model, we focus on two pubic and widely-used EHR datasets: MIMIC-III Johnson et al. (2016) and MIMIC-IV Johnson et al. (2020). We focus on MIMIC-III as the primary dataset for our experiment and divide it into 6000/500/1000 for the training, validation, and test sets, respectively. The basic statistics of MIMIC-III and MIMIC-IV and results of MIMIC-IV are shown as supplement at the Appendix E.
Note that, we selected MIMIC-III as the primary dataset for two key reasons: it facilitates direct comparison with existing baselines commonly evaluated on MIMIC-III, ensuring reproducibility, and it natively supports ICD-9 codes, aligning with knowledge encoding base in baselines and our Diagnosis KG. Using MIMIC-IV would require mapping from ICD-10, potentially introducing errors that could affect model evaluation.
@@ 4.1.2 Baselines
To evaluate the performance of our proposed model, we selected the following 10 EHR models as comparison methods: (1) RNN/CNN-based models: Dipole Ma et al. (2017), Deepr Nguyen et al. (2017), RETAIN Choi et al. (2016a) and Timeline Bai et al. (2018). (2) Graph-based models: GRAM Choi et al. (2017), KAME Ma et al. (2018), and CGL Lu et al. (2021). (3) Transformer-based models: G-BERT Shang et al. (2019), HiTANet Luo et al. (2020), and THAM Li et al. (2024).
Note that although GraphCare Jiang et al. (2023) is another KG-based method, it is not included as a baseline in experiments because its KGs cannot be directly transferred from CCS codes to ICD-9 codes due to limited GPU memory. Supportive details will be provided in the following case study. Implementation details are discussed in Appendix D.
@@
We conduct our experiments on two tasks as downstream examples for clinical prediction, following the similar settings as in previous studies Choi et al. (2017); Lu et al. (2021):

### v2

@@ 3.2 Generate Diagnosis KG
KG Normalization
Our approach for normalizing the combined KG $\mathcal{G}_{H}$ involves two key steps: (1) Uniform encoding system for entities: In addition to phenotype nodes sourced from HPO, we align disease and drug entities using the ICD-9-CM [10] and ATC [42] coding systems, both of which are widely used in EHR data. A cross-referencing method is employed to convert $\mathcal{G}_{M}$ and $\mathcal{G}_{N}$ into a shared hierarchy. Given that drugs are not the primary focus, we use ATC-4 to simplify the numerous synergistic interactions in DrugBank. (2) Clustering duplicated or unmatched triples: The LLM-prompting process is run multiple times, which can generate duplicated nodes and edges. Moreover, cross-referencing files may not cover all concepts, and unmatched entities leave noise in KG. To address these problems, we calculate cosine similarity for nodes and edges based on their embeddings. Using a predefined threshold $\theta$, we group entities within categories (i.e. “disease”, “drug”, “phenotype”, and “other”), and apply hierarchical clustering to reorganize edges.
Table 1: Statistical Comparison among KGs. (* means results might fluctuate by different parameter settings.)
@@
Advancement Analysis
To demonstrate the advancements of our proposed Diagnosis KG, we compare it with two recent health-related KGs: GraphCare [9] and PrimeKG [43]. A statistical summary of these KGs is presented in Table 1. The analysis highlights the following advantages of our KG:
@@
    Compared to GraphCare [9], our KG is larger in scale and is built upon the widely-used ICD-9-CM and HPO system, offering hierarchical and precise knowledge. In contrast, GraphCare relies on zero-shot generation from LLMs without fine-tuning, which can result in incorrect triples due to hallucination [12], along with unrelated entities (e.g. social impact) that are irrelevant to the clinical task. Furthermore, the subgraph sampling process can introduce instability into the output. Hence, our KG provides more stable and accurate results in downstream prediction tasks.
@@
    Compared to PrimeKG [43], our KG includes more edge types, primarily augmented by LLM-generated triples. The LLM reasoning process not only enhances connectivity within the KG but also converts unstructured text into structured triples, making information retrieval more efficient from single-modal data.
Although PrimeKG covers a broader range of medical concepts, it is more prone to data quality issues that can introduce bias into EHR predictions. For example, genes and proteins included in PrimeKG are seldom considered in physicians’ decision-making processes, particularly when it comes to diagnosis. In contrast, our diagnosis-driven KG contains numerous normalized triples with diverse edges, enhanced by LLMs’ in-context learning capabilities, making it highly reproducible and well-suited for various diagnosis-related healthcare predictions. Such analysis will then be supported by the results of the ablation study.
@@ 3.3 Polar-space KG Embedding
Traditional KG embedding methods often emphasize either contextual or hierarchical information, limiting their ability to represent both simultaneously. In our setting, the constructed Diagnosis KG encodes both types of signals, requiring an embedding method capable of capturing their interplay. To meet this need, we adopt polar-space embedding [44], which balances semantic expressiveness with computational efficiency and interpretability. A comparison with other KG embedding methods [45, 46, 47, 48] is provided in Appendix 5.1.
To describe medical concepts from both contextual and hierarchical perspectives, we leverage two properties of the polar coordinate system: (1) Radial coordinate: map entities across different levels of the hierarchy; (2) Angular coordinate: map entities at the same level based on contextual information. We denote entities, including head $\mathbf{h}$ and tail $\mathbf{t}$, as $\mathbf{e}$, and relation embeddings as $\mathbf{r}$. Subscripts, such as $\mathbf{e}_{r}$ and $\mathbf{e}_{a}$, represent the modulus and phase components, respectively. Each entry of $\mathbf{h}_{r,a}$ or $\mathbf{t}_{r,a}$ corresponds to a radial or angular value, while each entry of $\mathbf{r}_{r,a}$ represents a scaling transformation between the corresponding head and tail. The radial and angular coordinates can be formulated as follows:
@@
where $\|\cdot\|_{1}$, $\|\cdot\|_{2}$ denote the $\ell_{1}$, $\ell_{2}$ norm. The final circular arc distance $d(\mathbf{h},\mathbf{t})$ is the weighted sum of both radial and angular distances. Note that, we can get $\mathbf{h}$ and $\mathbf{t}$ by concatenating both radial and angular vectors. Such embedding model can then be pretrained individually through link prediction tasks, which helps us understand priors from KG into trainable embeddings. The regular negative sampling loss function $L$ is adopted for optimization:
@@
where $\gamma$ and $\sigma$ are marginal value and sigmoid function respectively, and $({h^{\prime}},r,{t^{\prime}})$ is a negative triple. Therefore, we can get diagnosis embedding matrix $\mathbf{E_{\mathcal{D}}}$ for condition codes.
Note that we also experiment with a simpler objective function (i.e., bringing connected nodes closer while pushing unconnected nodes apart) as a replacement for the link prediction task. However, due to the complexity of the KG, the model fails to fully learn the positions of nodes (refer to the ablation study in section 4.3).
3.4 Encode Graph-Enhanced Patient Embeddings
The continual enhancement process unfolds in two steps: graph learning and proxy-task learning, which together refine the priors from the “Knowledge Scholar” by fully leveraging EHR data. The proposed encoder-decoder framework not only emphasizes the comprehensive understanding of medical patterns but also seamlessly integrates local knowledge into the learning process, further capturing critical medical patterns for precise diagnosis. The backbone of the encoder module contains graph learning and attention, which adjusts code-level embeddings and progressively refines them into comprehensive patient representations.
@@ 3.4 Encode Graph-Enhanced Patient Embeddings
$\mathbf{W}_{c}$, $\mathbf{W}_{u}$, and $\mathbf{W}_{v}$ are weighted matrices. The attention score $\alpha_{\tau}$ measures the distribution of medical codes of admission, and $\beta$ measures the distribution across admissions. Note that, we project each admission $\mathbf{v}_{\tau}$ to $\tilde{\mathbf{v}}_{\tau}$ for fitting the patient dimension through a single layer with $\mathbf{W}_{u}$.
The proposed encoder module ensures that both the visit-level diagnostic information and the longitudinal history of a patient’s health are effectively captured in the final patient embedding $\mathbf{p}$, laying a robust foundation for the subsequent self-supervised learning phase.
3.5 Stepwise Lab-Informed Pretraining
Most existing methods [4, 5, 6] adapt pretraining tasks from language models to exploit the whole EHR dataset, but fewer of them consider whether definitions are clinically meaningful. Motivated by this gap, we aim to harness laboratory insights to help predictive models better mimic the clinical reasoning process of a human doctor. Hence, we introduce a novel proxy-task learning paradigm that combine two clinically grounded steps, (1) lab test assignment and (2) abnormality detection, into the pretraining phase, thereby aligning the training process more closely with stepwise diagnostic reasoning. To the best of our knowledge, this is the first attempt to explicitly incorporate lab test outcomes as labels during pretraining.
The EHR dataset for the proxy task learning is denoted as $\mathcal{S}^{\prime}$ with $N^{\prime}$ patients in total. Since those single-admission patients and final admission records of multi-admission patient lack labels regarding their next admission, we assume that the learned representations $\{\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{N^{\prime}}\}$ will be able to reflect their medical histories (i.e. historical lab test results). Considering diverse patterns across different lab tests within the vocabulary =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{|\mathcal{L}|}\}$, we further categorize lab tests into several distinct clusters. Note that, the categorical information of each test is documented in most EHR datasets. Here we take lab tests in MIMIC datasets as an example:
@@ 3.5 Stepwise Lab-Informed Pretraining
    Hematology (# items: 417) Analyze blood components like red and white blood cells to monitor conditions.
@@
However, it still remains challenging to enable models to grasp the diagnostic reasoning process. Typically, a doctor first forms preliminary diagnostic hypotheses based on symptoms and medical history, orders relevant lab tests accordingly, and then confirms the diagnosis based on abnormal test results. This sequential reasoning underscores how to quantify conditional probabilities, which assign lab tests before observing abnormal outcomes, for our proxy-task learning module.
For better demonstration, we then separate the vocabulary $\mathcal{L}$ into $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ in order and take single patient as an example. Since we aim to decode embedding $\mathbf{p}$ into the probability distribution $\mathbf{\hat{y}}:\mathbf{\hat{y}}_{i}=P(l_{i}\in\mathcal{L}_{m}|\mathbf{p})$ for each category $m$, it is intuitive that the equation can be parameterized by a multi-layer perceptron (MLP) for estimation $\mathbf{\hat{y}}=\sigma(\text{MLP}(\mathbf{p}))$. However, such naïve method has limitation in mimicking the conditional probability for both assignment or abnormality step. Assuming assigned and abnormal lab results are denoted as $\mathbf{\hat{y}}^{a,m}$ and $\mathbf{\hat{y}}^{b,m}$ respectively for certain category $m$, the abnormal lab results $\mathbf{\hat{y}}^{b,m}$ is equivalent to a joint probability $\text{P}\left(p^{a}_{l_{i}}\in\mathcal{L}_{m},p^{b}_{l_{i}}\in\mathcal{L}_{m}|\mathbf{p}\right)$ regarding assignment as priors, which can be represented as follow with given $\mathbf{p}$:
$\mathbf{\hat{y}}^{b,m}$ $=\text{P}\left(p^{a}_{l_{i}}\in\mathcal{L}_{m},p^{b}_{l_{i}}\in\mathcal{L}_{m}|\mathbf{p}\right)$
$=\text{P}\left(p^{b}_{l_{i}}\in\mathcal{L}_{m}|p^{a}_{l_{i}}\in\mathcal{L}_{m},\mathbf{p}\right)\times\text{P}\left(p^{a}_{l_{i}}\in\mathcal{L}_{m}|\mathbf{p}\right)$ (12)
where $i$ represent the index within the vocabulary $\mathcal{L}$, so that $p^{a}_{l_{i}}$ and $p^{b}_{l_{i}}$ are probabilities of the $i$-th assigned and abnormal lab results. Even though we separate our target $\mathbf{\hat{y}}^{b}$ by two consecutive conditional components, it is still difficult to parameterize $\text{P}\left(p^{b}_{l_{i}}\in\mathcal{L}_{m}|p^{a}_{l_{i}}\in\mathcal{L}_{m},\mathbf{p}\right)$ for estimation. Hence, we adopt the constraint relaxation method and parameterize $\text{P}\left(p^{b}_{l_{i}}\in\mathcal{L}_{m}|\mathbf{p}\right)$ as the lower bound approximation:
$\text{P}\left(p^{b}_{l_{i}}\in\mathcal{L}_{m}|p^{a}_{l_{i}}\in\mathcal{L}_{m},\mathbf{p}\right)$ $\geq\text{P}\left(p^{b}_{l_{i}}\in\mathcal{L}_{m}|\mathbf{p}\right)$ (13)
$\text{P}\left(p^{a}_{l_{i}}\in\mathcal{L}_{m}|\mathbf{p}\right)$ $=\sigma(\mathbf{w}_{a}\mathbf{p})_{l_{i}}$ (14)
$\text{P}\left(p^{b}_{l_{i}}\in\mathcal{L}_{m}|\mathbf{p}\right)$ $=\sigma(\mathbf{w}_{b}\mathbf{p})_{l_{i}}$ (15)
$\mathbf{\hat{y}}^{b,m}_{i}$ $=[\sigma(\mathbf{w}^{m}_{a}\mathbf{p})\otimes\sigma(\mathbf{w}^{m}_{b}\mathbf{p})]_{l_{i}}$ (16)
$L^{m}=\;$ $\omega\times\text{BCE}(\mathbf{\hat{y}}^{a,m},\mathbf{y}^{a,m})$ (17)
$+(1-\omega)\times\text{BCE}(\mathbf{\hat{y}}^{b,m},\mathbf{y}^{b,m})$
where $\mathbf{w}^{m}_{a},\mathbf{w}^{m}_{b}\in\mathbb{R}^{|\mathcal{L}_{m}|\times p}$ decodes patient embedding $\mathbf{p}$ to the probability of the laboratory codes for assignment and abnormality respectively. In $L^{m}$, $\text{BCE}(\cdot)$ denotes the binary cross-entropy loss and $\omega$ is the hyperparameter scaling two losses. $\mathbf{y}^{a,m}$ and $\mathbf{y}^{b,m}$ are ground truth for prediction. Note that, as a self-supervised learning problem, such design help model learn the conditional relationship between both steps, which further augment the robustness of learned parameters. Moreover, to ensure the lab embeddings are well-aligned with each category, the proxy task will be trained in two phrases:
  • Joint Training. We get preliminary trained parameters in encoder $\Theta_{e}$ and optimize prediction through the average loss across three categories, where we can adjust patient embedding $\tilde{\mathbf{p}}$ through updated parameters.
@@ 4.1.1 Datasets
To evaluate our proposed model, we utilize two public and widely-used EHR datasets: MIMIC-III [49] and MIMIC-IV [50]. MIMIC-III includes 7,493 patients with multiple visits ($T\geq 2$) recorded between 2001 and 2012, while MIMIC-IV contains 85,155 such patients from 2008 to 2019. To ensure a comparable sample size and avoid temporal overlap with MIMIC-III, we randomly sample 10,000 patients from MIMIC-IV who were admitted between 2013 and 2019. Table 2 summarizes the basic statistics of the MIMIC-III and MIMIC-IV datasets. We divide both datasets into training, validation, and test sets by patient: 6,000/493/1,000 patients for MIMIC-III and 8,000/1,000/1,000 for MIMIC-IV, respectively. For each patient, the most recent visit is used as the prediction target, with all previous visits serving as input features. All diagnoses are encoded using the ICD-9-CM coding system.
Table 2: Statistics of MIMIC-III and MIMIC-IV datasets
Dataset MIMIC-III MIMIC-IV
# patients 7,493 10,000
Max. # visit 42 55
Avg. # visit 2.66 3.66
# codes 4,880 6,102
Max. # codes per visit 39 50
Avg. # codes per visit 13.06 13.38
@@ 4.1.2 Baselines
To evaluate the performance of our proposed model, we selected the following 10 EHR models as comparison methods: (1) RNN/CNN-based models: RETAIN [17] and Timeline [20]. (2) Graph/KG-based models: GCT [2], CGL [1], SeqCare [11], and RAM-EHR [12]. (3) Transformer-based models: G-BERT [4], HiTANet [51], Med-BERT [24], and GT-BEHRT [6].
Note that although GraphCare [9] is another KG-based method, it is not included as a baseline in experiments because its defined KGs cannot be directly transferred from CCS codes to ICD-9 codes. However, we feed our diagnosis KG into GraphCare framework for comparison as shown in Table 5.
@@
We conduct our experiments on two tasks as downstream examples for clinical prediction, following the similar settings as in previous studies [7, 1]:
@@
4.1.4 Implementation Details
During the transformation of textual information into triples via LLMs, each inference undergoes iterative prompting and re-read prompting strategies with two and one rounds, respectively. Polar-space hierarchical KG embeddings are used as node features in the graph learning module, with vector size $20002k=$ 2000 ($k=1000$ for both hierarchical and semantic components). The disease-lab graph is modeled using two GAT convolution layers with hidden dimensions of 256, and default settings elsewhere. Attention modules adopt dimensions $a=256$ and $b=256$, with a dropout rate of 0.2 applied to each. Lab test embeddings are generated using three decoders, each implemented as a three-layer MLP with hidden sizes of 256 and 128, and a uniform dropout rate of 0.4 across all layers for both pretraining and direct training. A two-layer classifier with 256 hidden units and a dropout rate of 0.5 is used to transform concatenated features into logits for downstream tasks. KG embeddings are trained for 180,000 steps to capture hierarchical and semantic relations. Proxy-task learning is performed with 10 epochs for both joint and individual training. Diagnosis and heart failure predictions use 500 and 50 training epochs, respectively. The Adam optimizer with a decaying learning rate is employed throughout, with an initial learning rate of 0.001 for both pretraining and finetuning. All experiments are conducted using Python 3.10 and PyTorch 2.3.1 with CUDA 12.4 on a system equipped with two AMD EPYC 9254 24-Core processors, 528 GB RAM, and four NVIDIA L40S GPUs.